# Composite Learning Backstepping Control With Guaranteed Exponential Stability and Robustness

### v3

@@ Appendix C The proof of Theorem 3
$-\kappa(1+p)\tilde{\bm{\theta}}^{T}(Q(t,t_{\rm e}){\tilde{\bm{\theta}}}+{\bm{d}}_{\rm g}).$
@@
$\dot{V}\leq$ $-\bm{e}^{T}(K-I/4)\bm{e}-\kappa\tilde{\bm{\theta}}^{T}Q(t,t_{\rm e}){\tilde{\bm{\theta}}}+\bm{e}^{T}(\Phi-\Phi_{\rm f})\tilde{\bm{\theta}}$
$+\bm{e}^{T}\bm{d}-(1+p)\tilde{\bm{\theta}}^{T}\Phi_{{\rm f}}{\bm{d}}_{\rm f}-\kappa(1+p)\tilde{\bm{\theta}}^{T}{\bm{d}}_{\rm g}.$
@@
$+\bm{e}^{T}\bm{d}-\kappa(1+p)\tilde{\bm{\theta}}^{T}{\bm{d}}_{\rm g}$
@@
$\dot{V}\leq$ $-k_{\rm c}\|\bm{e}\|^{2}+\delta c_{\theta}\|\bm{e}\|+\bar{d}\|\bm{e}\|+(1+p)(\bar{\epsilon}\bar{d}+\kappa c_{\theta}{\bar{d}}_{\rm g}).$
@@
$\dot{V}\leq-k_{\rm c}\|\bm{e}\|^{2}/2+\rho(k_{\rm c},\kappa,\bar{d})$
with $\rho$ :=:= $(\delta c_{\theta}+\bar{d})^{2}/(2k_{\rm c})+(1+p)(\bar{\epsilon}\bar{d}+\kappa c_{\theta}{\bar{d}}_{\rm g})\in\mathbb{R}^{+}$, which is valid on $\bm{x}\in\Omega_{{\rm c}_{x}}$ for $t\in[0,t_{\rm f})$. The remaining derivation is similar to the steps from (44) to the end of the proof of Item 1 in Theorem 2, so it is omitted here for saving space.
@@
${\dot{V}}_{\zeta}=$ $\bm{e}^{T}(-K\bm{e}+\Phi^{T}_{\zeta}{\tilde{\bm{\theta}}}_{\zeta})+\bm{e}^{T}\bm{d}-(1+p){\tilde{\bm{\theta}}}_{\zeta}^{T}\Phi_{{\rm f},\zeta}(\Phi_{{\rm f},\zeta}^{T}\tilde{\bm{\theta}}_{\zeta}+{\bm{d}}_{\rm f})$
$-\kappa(1+p)\tilde{\bm{\theta}}^{T}_{\zeta}(Q_{\zeta}(t,t_{\rm e}){\tilde{\bm{\theta}}}_{\zeta}+{\bm{d}}_{\rm g}).$
@@
$\dot{V}_{\zeta}\leq$ $-k_{\rm c}\bm{e}^{T}\bm{e}-\kappa(1+p)\sigma^{*}\tilde{\bm{\theta}}^{T}_{\zeta}\tilde{\bm{\theta}}_{\zeta}+\delta\|\bm{e}\|\|\tilde{\bm{\theta}}_{\zeta}\|$
$+\bm{e}^{T}\bm{d}-(1+p){\tilde{\bm{\theta}}}_{\zeta}^{T}\Phi_{{\rm f},\zeta}{\bm{d}}_{\rm f}-\kappa(1+p){\tilde{\bm{\theta}}}_{\zeta}^{T}{\bm{d}}_{\rm g}.$
@@
$\dot{V}_{\zeta}\leq$ $-k_{\rm c}\|\bm{e}\|^{2}-\kappa(1+p)\sigma^{*}\|\tilde{\bm{\theta}}_{\zeta}\|^{2}+\delta\|\bm{e}\|\|\tilde{\bm{\theta}}_{\zeta}\|$
$+\bar{d}\|\bm{e}\|+(1+p)(\bar{\epsilon}\bar{d}+\kappa{\bar{d}}_{\rm g}\|\tilde{\bm{\theta}}_{\zeta}\|).$
@@
$\dot{V}_{\zeta}\leq$ $-k_{\rm c}\|\bm{e}\|^{2}/2+\bar{d}\|\bm{e}\|-\kappa\sigma^{*}\|\tilde{\bm{\theta}}_{\zeta}\|^{2}$
$+(1+p)\kappa{\bar{d}}_{\rm g}\|\tilde{\bm{\theta}}_{\zeta}\|+(1+p)\bar{\epsilon}\bar{d}$
with $p$ == $\delta^{2}/(2k_{\rm c}\kappa\sigma^{*})\in\mathbb{R}^{+}$, which can be rewritten into
@@
$-\kappa\sigma^{*}\mu(\|\tilde{\bm{\theta}}_{\zeta}\|^{2}-(1+p){\bar{d}}_{\rm g}/(\sigma^{*}\mu)\|\tilde{\bm{\theta}}_{\zeta}\|)$
$-\kappa\sigma^{*}(1-\mu)\|\tilde{\bm{\theta}}_{\zeta}\|^{2}+(1+p)\bar{\epsilon}\bar{d}$
@@
$\dot{V}_{\zeta}\leq$ $-k_{\rm c}(1-\mu)\|\bm{e}\|^{2}/2-\kappa\sigma^{*}(1-\mu)\|\tilde{\bm{\theta}}_{\zeta}\|^{2}+\rho_{2}(k_{\rm c},\kappa,\bar{d})$
with $\rho_{2}(k_{\rm c},\kappa,\bar{d})$ :=:= $(1+p)\bar{\epsilon}\bar{d}+\bar{d}^{2}/(2k_{\rm c}\mu)+\kappa((1+p){\bar{d}}_{\rm g})^{2}$ $(4\sigma^{*}\mu) \in$ $\mathbb{R}^{+}$. It follows from the above result that
$\dot{V}_{\zeta}\leq$ $-k_{\rm c}(1-\mu)\bm{e}^{T}\bm{e}/2-\kappa\sigma^{*}\lambda_{\min}(\Gamma_{\zeta})(1-\mu)\tilde{\bm{\theta}}_{\zeta}^{T}\Gamma^{-1}_{\zeta}\tilde{\bm{\theta}}_{\zeta}$
$+\rho_{2}(k_{\rm c},\kappa,\bar{d}).$
@@
$\dot{V}_{\zeta}\leq$ $-k_{\rm r}V_{\zeta}+\rho_{2}(k_{\rm c},\kappa,\bar{d}),\forall t\geq T_{\rm a}$
with $k_{\rm r}$ :=:= $\min\{k_{\rm c}(1-\mu),2\kappa\sigma^{*}\lambda_{\min}(\Gamma_{\zeta})(1-\mu)/(1+p)\}\in\mathbb{R}^{+}$. Applying [88, Lemma A.3.2] to the above inequality yields
$V_{\zeta}(t)\leq(V_{\zeta}(0)-\rho_{2}(k_{\rm c},\kappa,\bar{d}))e^{-k_{\rm r}t}+\rho_{2}(k_{\rm c},\kappa,\bar{d}),\forall t\geq T_{\rm a}$
which impiles that the equilibrium point $(\bm{e}$, $\tilde{\bm{\theta}})=\bm{0}$ of the closed-loop system (25) with (23) has partial practical exponential stability on $t$ $\in$ $T_{\rm a},\infty)$, where the tracking error $\bm{e}$ and the partial estimation error $\tilde{\bm{\theta}}_{\zeta}$ exponentially converge to a small neighborhood of $\bm{0}$ dominated by $k_{{\rm c}i}$, $\kappa$, and $\bar{d}$.
@@
$\dot{V}\leq$ $-k_{\rm c}(1-\mu)\bm{e}^{T}\bm{e}/2-\kappa\sigma^{*}\lambda_{\min}(\Gamma)(1-\mu)\tilde{\bm{\theta}}^{T}\Gamma^{-1}\tilde{\bm{\theta}}+\rho_{2}$
$\leq$ $-k_{\rm r}V+\rho_{2}(k_{\rm c},\kappa,\bar{d}),\forall t\geq T_{\rm e}$
with $k_{\rm r}$ :=:= $\min\{k_{\rm c}(1-\mu),2\kappa\sigma^{*}\lambda_{\min}(\Gamma)(1-\mu)/(1+p)\}\in\mathbb{R}^{+}$, $\rho_{2}$ :=:= $(1+p)\bar{\epsilon}\bar{d}+\bar{d}^{2}/(2k_{\rm c}\mu)+\kappa((1+p){\bar{d}}_{\rm g})^{2}/(4\sigma^{*}\mu)\in\mathbb{R}^{+}$, and $p$ == $\delta^{2}/(2k_{\rm c}\kappa\sigma^{*})\in\mathbb{R}^{+}$, where $\delta\in\mathbb{R}^{+}$ satisfies $\|\Phi-\Phi_{\rm f}\|$ $\leq$ $\delta$, $\forall t\geq 0$, and $\sigma^{*}\in\mathbb{R}^{+}$ with $\sigma^{*}\leq\sigma$ satisfies (30). Thus, the equilibrium point $(\bm{e}$, $\tilde{\bm{\theta}})=\bm{0}$ of the closed-loop system (25) with (23) is practical exponential stable on $t$ $\in$ $T_{\rm e},\infty)$, where the tracking error $\bm{e}$ and the estimation error $\tilde{\bm{\theta}}$ exponentially converge to a small neighborhood of $\bm{0}$ dominated by $k_{{\rm c}i}$, $\kappa$, and $\bar{d}$. ∎

### v4

@@ Appendix C The proof of Theorem 3
$-\kappa_{2}(1+p)\tilde{\bm{\theta}}^{T}(Q(t,t_{\rm e}){\tilde{\bm{\theta}}}+{\bm{d}}_{\rm g}).$
@@
$\dot{V}\leq$ $-\bm{e}^{T}(K-I/4)\bm{e}-\kappa_{2}\tilde{\bm{\theta}}^{T}Q(t,t_{\rm e}){\tilde{\bm{\theta}}}+\bm{e}^{T}(\Phi-\Phi_{\rm f})\tilde{\bm{\theta}}$
$+\bm{e}^{T}\bm{d}-(1+p)\tilde{\bm{\theta}}^{T}\Phi_{{\rm f}}{\bm{d}}_{\rm f}-\kappa_{2}(1+p)\tilde{\bm{\theta}}^{T}{\bm{d}}_{\rm g}.$
@@
$+\bm{e}^{T}\bm{d}-\kappa_{2}(1+p)\tilde{\bm{\theta}}^{T}{\bm{d}}_{\rm g}$
@@
$\dot{V}\leq$ $-k_{\rm c}\|\bm{e}\|^{2}+\delta c_{\theta}\|\bm{e}\|+\bar{d}\|\bm{e}\|+(1+p)(\bar{\epsilon}\bar{d}+\kappa_{2}c_{\theta}{\bar{d}}_{\rm g}).$
@@
$\dot{V}\leq-k_{\rm c}\|\bm{e}\|^{2}/2+\rho(k_{\rm c},\kappa_{2},\bar{d})$
with $\rho$ :=:= $(\delta c_{\theta}+\bar{d})^{2}/(2k_{\rm c})+(1+p)(\bar{\epsilon}\bar{d}+\kappa_{2}c_{\theta}{\bar{d}}_{\rm g})\in\mathbb{R}^{+}$, which is valid on $\bm{x}\in\Omega_{{\rm c}_{x}}$ for $t\in[0,t_{\rm f})$. The remaining derivation is similar to the steps from (44) to the end of the proof of Item 1 in Theorem 2, so it is omitted here for saving space.
@@
${\dot{V}}_{\zeta}=$ $\bm{e}^{T}(-K\bm{e}+\Phi^{T}_{\zeta}{\tilde{\bm{\theta}}}_{\zeta})-(1+p){\tilde{\bm{\theta}}}_{\zeta}^{T}\Phi_{{\rm f},\zeta}(\Phi_{{\rm f},\zeta}^{T}\tilde{\bm{\theta}}_{\zeta}+{\bm{d}}_{\rm f})$
$-\kappa_{2}(1+p)\tilde{\bm{\theta}}^{T}_{\zeta}(Q_{\zeta}(t,t_{\rm e}){\tilde{\bm{\theta}}}_{\zeta}+{\bm{d}}_{\rm g})+\bm{e}^{T}\bm{d}.$
@@
$\dot{V}_{\zeta}\leq$ $-k_{\rm c}\bm{e}^{T}\bm{e}-\kappa_{2}(1+p)\sigma^{*}\tilde{\bm{\theta}}^{T}_{\zeta}\tilde{\bm{\theta}}_{\zeta}+\delta\|\bm{e}\|\|\tilde{\bm{\theta}}_{\zeta}\|$
$+\bm{e}^{T}\bm{d}-(1+p){\tilde{\bm{\theta}}}_{\zeta}^{T}\Phi_{{\rm f},\zeta}{\bm{d}}_{\rm f}-\kappa_{2}(1+p){\tilde{\bm{\theta}}}_{\zeta}^{T}{\bm{d}}_{\rm g}.$
@@
$\dot{V}_{\zeta}\leq$ $-k_{\rm c}\|\bm{e}\|^{2}-\kappa_{2}(1+p)\sigma^{*}\|\tilde{\bm{\theta}}_{\zeta}\|^{2}+\delta\|\bm{e}\|\|\tilde{\bm{\theta}}_{\zeta}\|$
$+\bar{d}\|\bm{e}\|+(1+p)(\bar{\epsilon}\bar{d}+\kappa_{2}{\bar{d}}_{\rm g}\|\tilde{\bm{\theta}}_{\zeta}\|).$
@@
$\dot{V}_{\zeta}\leq$ $-k_{\rm c}\|\bm{e}\|^{2}/2+\bar{d}\|\bm{e}\|-\kappa_{2}\sigma^{*}\|\tilde{\bm{\theta}}_{\zeta}\|^{2}$
$+(1+p)\kappa_{2}{\bar{d}}_{\rm g}\|\tilde{\bm{\theta}}_{\zeta}\|+(1+p)\bar{\epsilon}\bar{d}$
with $p$ == $\delta^{2}/(2k_{\rm c}\kappa_{2}\sigma^{*})\in\mathbb{R}^{+}$, which can be rewritten into
@@
$-\kappa_{2}\sigma^{*}\mu(\|\tilde{\bm{\theta}}_{\zeta}\|^{2}-(1+p){\bar{d}}_{\rm g}/(\sigma^{*}\mu)\|\tilde{\bm{\theta}}_{\zeta}\|)$
$-\kappa_{2}\sigma^{*}(1-\mu)\|\tilde{\bm{\theta}}_{\zeta}\|^{2}+(1+p)\bar{\epsilon}\bar{d}$
@@
$\dot{V}_{\zeta}\leq$ $-k_{\rm c}(1-\mu)\|\bm{e}\|^{2}/2-\kappa_{2}\sigma^{*}(1-\mu)\|\tilde{\bm{\theta}}_{\zeta}\|^{2}+\rho_{2}(k_{\rm c},\kappa_{2},\bar{d})$
with $\rho_{2}(k_{\rm c},\kappa_{2},\bar{d})$ :=:= $(1+p)\bar{\epsilon}\bar{d}+\bar{d}^{2}/(2k_{\rm c}\mu)+\kappa_{2}((1+p){\bar{d}}_{\rm g})^{2}$ $(4\sigma^{*}\mu) \in$ $\mathbb{R}^{+}$. It follows from the above result that
$\dot{V}_{\zeta}\leq$ $-k_{\rm c}(1-\mu)\bm{e}^{T}\bm{e}/2-\kappa_{2}\sigma^{*}(1-\mu)\tilde{\bm{\theta}}_{\zeta}^{T}\tilde{\bm{\theta}}_{\zeta}+\rho_{2}(k_{\rm c},\kappa_{2},\bar{d}).$
@@
$\dot{V}_{\zeta}\leq$ $-k_{\rm r}V_{\zeta}+\rho_{2}(k_{\rm c},\kappa_{2},\bar{d}),\forall t\geq T_{\rm a}$
with $k_{\rm r}$ :=:= $\min\{k_{\rm c}(1-\mu),2\kappa_{2}\sigma^{*}(1-\mu)/(1+p)\}\in\mathbb{R}^{+}$. Applying [42, Lemma A.3.2] to the above inequality yields
$V_{\zeta}(t)\leq(V_{\zeta}(0)-\rho_{2}(k_{\rm c},\kappa_{2},\bar{d}))e^{-k_{\rm r}t}+\rho_{2}(k_{\rm c},\kappa_{2},\bar{d}),\forall t\geq T_{\rm a}$
which implies that the equilibrium point $(\bm{e}$, $\tilde{\bm{\theta}})=\bm{0}$ of the closed-loop system (25) with (23) achieves partial practical exponential stability on $t$ $\in$ $T_{\rm a},\infty)$, where the tracking error $\bm{e}$ and the partial estimation error $\tilde{\bm{\theta}}_{\zeta}$ exponentially converge to a small neighborhood of $\bm{0}$ dominated by $k_{{\rm c}i}$, $\kappa_{1}$, $\kappa_{2}$, and $\bar{d}$.
@@
$\dot{V}\leq$ $-k_{\rm c}(1-\mu)\bm{e}^{T}\bm{e}/2-\kappa_{2}\sigma^{*}(1-\mu)\tilde{\bm{\theta}}^{T}\tilde{\bm{\theta}}+\rho_{2}$
$\leq$ $-k_{\rm r}V+\rho_{2}(k_{\rm c},\kappa_{2},\bar{d}),\forall t\geq T_{\rm e}$
with $k_{\rm r}$ :=:= $\min\{k_{\rm c}(1-\mu),2\kappa_{2}\sigma^{*}(1-\mu)/(1+p)\}\in\mathbb{R}^{+}$, $\rho_{2}$ :=:= $(1+p)\bar{\epsilon}\bar{d}+\bar{d}^{2}/(2k_{\rm c}\mu)+\kappa_{2}((1+p){\bar{d}}_{\rm g})^{2}/(4\sigma^{*}\mu)\in\mathbb{R}^{+}$, and $p$ == $\delta^{2}/(2k_{\rm c}\kappa_{2}\sigma^{*})\in\mathbb{R}^{+}$, where $\delta\in\mathbb{R}^{+}$ satisfies $\|\Phi-\Phi_{\rm f}\|$ $\leq$ $\delta$, $\forall t\geq 0$, and $\sigma^{*}\in\mathbb{R}^{+}$ ($\sigma^{*}\leq\sigma$) satisfies (30). Consequently, the equilibrium point $(\bm{e}$, $\tilde{\bm{\theta}})=\bm{0}$ of the closed-loop system (25) with (23) achieves practical exponential stability on $t$ $\in$ $T_{\rm e},\infty)$, where the tracking error $\bm{e}$ and the estimation error $\tilde{\bm{\theta}}$ converge to a small neighborhood of $\bm{0}$ dominated by $k_{{\rm c}i}$, $\kappa_{1}$, $\kappa_{2}$, and $\bar{d}$ exponentially. ∎